You are here: Foswiki>Commissioning Web>Logs>24Aug013_Log (14 Sep 2013, JuanGuerra)Edit Attach

UT 20130213 AO#2 SX

Telescope Operator: David(LBTO)
SW Support:
AO Support:
Telescope Support:
Instrument Support: John M (LBTO)


  • Summer shutdown.
  • The DX Unit is back on the telescope with the spare thin shell TS4.
  • Two Arcetri team in the telescope; one to work on ASM and another one to work in the Py-WFS stages for DX and SX.
  • 4D+ Retroreflector aligned two days before by Lee and Guido.
  • Cooling system up and running.
  • First check of the 4D intereferometer with the ASM_DX. It has been found some problems to get the whole pupil.


The 4D was aligned some days before. The alignment was good and the first task working with the 4D is to get the full image pupil in the interferogram. During the check process of the 4D+ASM it was found a problem to determine if the trigger was working properly. The trigger fiber was checked in the upper tree house and in the lower tree house checking the output signal with a voltimeter connected to the trigger box. (include some photos).

Once the trigger was running perfectly the Arcetri team decided to rest till midnight and to work the rest of the night. Runa and Juan Carlos keep working in the alignment of the pupil.

We have crossed problem to get a full image of the pupil, the OD is getting hard to flat.

The process of Gaussian or Modes applied(0 to 30) help a bit to have a better shape, but we are still having problem with most of the OD.

3:00 (LT) Exploring new procedure to apply to the mirror some local shape.

-- JuanGuerra - 25 Aug 2013
  • m3 and m2 positions for the 4D:
Topic revision: r3 - 14 Sep 2013, JuanGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback